
Geo Semester B Unit 6 Notes
Presentation
•
Mathematics
•
9th - 12th Grade
•
Easy
Victoria Colbert
Used 1+ times
FREE Resource
20 Slides • 21 Questions
1
2
3
4
Multiple Choice
What is the volume of the prism given?
72 cm3
144 cm3
288 cm3
432 cm3
5
6
Multiple Choice
What is the volume of the cylinder given?
62.8 cm2
62.8 cm3
125.6 cm3
125.6 in3
7
8
Multiple Choice
What is the volume of the pyramid?
432 cm2
144 cm2
432 cm3
144 cm3
9
Multiple Choice
What is the volume of a rectangular pyramid with the dimension of the base 5 in x 3 in, and height 8 in?
20 in3
40 in3
80 in3
120 in3
10
11
Multiple Choice
What is the volume of the cone?
103.62 cm3
310.86 cm3
932.58 cm3
932.58 in3
12
Multiple Choice
What is the volume of the cone?
103.62 cm3
310.86 cm3
932.58 cm3
932.58 in3
13
14
Multiple Choice
What is the volume of the sphere?
100.2 cm3
276.55 cm3
267.95 cm3
803.84 cm3
15
Multiple Choice
What is the volume of the sphere with a diameter of 15 m?
1592.55 m3
1766.25 m3
5295.75 m3
6005.75 m3
16
​
17
18
Some of the coins are larger than the others. There is no easy way to tell whether the stacks have the same volume.
19

20
Independent - One activity does not effect the outcome of a different activity. (Drawing marbles from a bag, but putting them back each time.)
Dependent - One activity DOES effect the outcome of another activity. (Drawing a card from a deck of cards, not putting it back and drawing another.)
21
Multiple Select
You roll a number cube twice. The first time is a 3 and the second time is an even number.
Independent
Dependent
22
Multiple Select
You randomly draw a marble from a bag of marbles. You get a red and don't put it back into the bag. You draw another marble out of the bag and get a yellow.
Independent
Dependent
23
Use the formula for the probability of independent events.
P(A and B) = P(A)⋅P(B)
Example: Find the probability of spinning an odd number on a spinner numbered 1 - 5; and the probability of flipping a coin and landing on tails.
24
Use the formula for the probability of independent events.
P(A and B) = P(A)⋅P(B)
Example: Find the probability of spinning an odd number on a spinner numbered 1 - 5; and the probability of flipping a coin and landing on tails.
25
Multiple Choice
Find the probability of spinning a 2 on a spinner numbered 1 - 5; and the probability of flipping a coin and landing on tails.
1/10 or 10%
1/4 or 25%
1/2 or 50%
2/3 or 67%
26
Multiple Choice
People are randomly chosen to be game show contestants from an audience of 100 people. You are with 5 of your relatives and 6 other friends. What is the probability that you, your relatives, and your friends are not chosen to be either of the first two contestants?
50/70 or 71.4%
58/75 or 77.3%
60/70 or 85.7%
88/100 or 88%
27
You roll a number cube and flip a coin. What is the probability of rolling a number greater than 4 and flipping tails? Use a table to find the sample space.
P(greater than 4 and tails)
28
Fill in the Blanks
What is the probability of rolling at most 4 and flipping heads?
Type answer...
29
When two events CANNOT happen at the same time, the events are said to be MUTUALLY EXCLUSIVE.
An example of this is getting heads on a coin and a tail on the same coin in the same toss.
Another example would be getting a six on a regular die and a five on the same die in the same roll.
A pre-Covid example would be being at school and being at home at the same time.
30
Multiple Choice
If you are picking a card randomly from a deck of cards, the events of picking a jack and picking a heart are ...
Mutually Exclusive
Not Mutually Exclusive
31
Multiple Choice
If you are picking a card randomly from a deck of cards, the events of picking an ace and picking a ‘3’ are ...
Mutually Exclusive
Not Mutually Exclusive
32
Example: Arif throws a biased coin. The probability of getting tails is 0.7. Therefore, the probability of getting heads is 1 - 0.7 = 0.3
33
Multiple Choice
Said throws a biased coin. The probability of getting tails is 0.4.
Work out the probability of getting heads.
0.4
0.6
0.8
0.2
34
To the right you see cards that can each be described in two ways
Jack, Queen, King or Ace
Heart, Club, Diamond or Spade
So, one card can be two things at once, for example a Queen and a Heart.
35
When we talk about OR probabilities we have to take "double identity" into account.
P(Queen or Heart) seems simple... just add the Queen probability to the Heart probability.... BUT
What about the Queen of Hearts? It will get counted twice, so....
We will have to consider OR and take away AND to get rid of the double count
36
P(A or B) = P(A) + P(B) - P(A and B)
P(Queen OR Heart) = P(Queen) + P(Heart) - P(Queen of Hearts)
= 4/16 + 4/16 - 1/16
P(Queen or Heart) = 7/16
37
Multiple Choice
38
Multiple Choice
39
Multiple Choice
If you roll one die, what is the probability of getting an even number or a multiple of 3?
(Looking at the picture might help you think about this)
1/3
2/3
1/2
1/6
40
Multiple Choice
If you roll one die, what is the probability of getting an even number or a multiple of 3?
(Looking at the picture might help you think about this)
1/3
2/3
1/2
1/6
41
Multiple Choice
The enrollment at Southburg High School is 1400. Suppose 550 students take French, 700 take algebra, and 400 take both French and algebra. What is the probability that a student selected randomly takes French or algebra?
Hint: P(French) + P(Algebra) - P(French and Algebra)
1250/1400
700/1400
550/1400
17/28
Show answer
Auto Play
Slide 1 / 41
SLIDE
Similar Resources on Wayground
37 questions
Geometry Section 11.1 Circles Pt 1
Presentation
•
9th - 12th Grade
37 questions
Circle Equations
Presentation
•
9th - 12th Grade
36 questions
Unit Circle Trig Practice
Presentation
•
9th - 12th Grade
33 questions
Amortization
Presentation
•
9th - 12th Grade
36 questions
Conjugaciones -ar, -er, -ir
Presentation
•
9th - 12th Grade
35 questions
Parts of the Parabola ALG 1
Presentation
•
9th - 12th Grade
35 questions
Biomolecule terms Part 1
Presentation
•
9th - 12th Grade
35 questions
The Scientific Method Notes Check
Presentation
•
9th - 12th Grade
Popular Resources on Wayground
10 questions
Safety Day 2026 – Student Review
Quiz
•
9th Grade
21 questions
Continents and Oceans
Quiz
•
6th Grade
20 questions
Parts of Speech
Quiz
•
5th Grade
16 questions
Subject & Predicate
Quiz
•
5th Grade
25 questions
Multiplication Facts
Quiz
•
5th Grade
12 questions
Map Skills
Quiz
•
3rd Grade
22 questions
Continents and Oceans
Quiz
•
5th Grade
15 questions
Rounding Decimals
Quiz
•
5th Grade
Discover more resources for Mathematics
12 questions
Absolute Value Equations
Quiz
•
9th Grade
20 questions
Function or Not a Function
Quiz
•
8th - 9th Grade
11 questions
Slope from 2 points
Quiz
•
8th - 9th Grade
10 questions
Exploring the Basics of Exponents
Interactive video
•
6th - 10th Grade
10 questions
Solving Absolute Value Equations
Quiz
•
9th Grade
20 questions
Unit 5 Test RSC (25-26)
Quiz
•
10th Grade
18 questions
SB 12.2 Write in point slope form
Quiz
•
9th Grade
12 questions
Identify Rigid Transformations
Quiz
•
8th - 12th Grade